40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m elog  Not logged in ELOG logo
Entry  Tue Mar 24 05:05:09 2015, ericq, Update, LSC, AO Path engaged CARM_DARM_AOengaged_23March2015.pdfloops.png
    Reply  Tue Mar 24 18:22:11 2015, ericq, Update, LSC, AO Path engaged IMC_ao_Mar242015.pngIMC_olgs_Mar242015.png
    Reply  Tue Mar 24 18:47:10 2015, ericq, Update, LSC, AO Path engaged 
Message ID: 11163     Entry time: Tue Mar 24 05:05:09 2015     Reply to this: 11167   11168
Author: ericq 
Type: Update 
Category: LSC 
Subject: AO Path engaged 

[J, Q]

Terse tonight, more verbose tomorrow. 

We have succesfully achieved multiple kHz bandwidth using the CARM AO path. The CM board super boosts are at too high of a frequency to use effectively, given the flattening of the AO TF. 


Jenne's totally, completely, and in all possible ways uncalibrated plot.  Calibration lines are in here (numbers in control room notebook).  I'm going to export and replot the data tomorrow, in real units.

CARM_DARM_AOengaged_23March2015.pdf

Attachment 1: CARM_DARM_AOengaged_23March2015.pdf  121 kB  Uploaded Tue Mar 24 06:06:03 2015  | Hide | Hide all | Show all
CARM_DARM_AOengaged_23March2015.pdf
Attachment 2: loops.png  417 kB  Uploaded Tue Mar 24 06:13:23 2015  | Show | Hide all | Show all
ELOG V3.1.3-